# Towards Million-Scale Adversarial Robustness Evaluation With Stronger Individual Attacks

### v2

@@ 2 Related Work
The accuracy of robustness evaluations relies heavily on the strength of the attack. White-box attacks used for robustness evaluation must be resilient to issues such as obfuscated gradients [6] and imbalanced gradients [7]. As shown in [6], standard PGD can overestimate robustness in the presence of obfuscated gradients, leading to the development of methods aimed at improving PGD’s effectiveness. These improvements can be broadly classified into two categories: 1) attack strategies and 2) loss functions. Improved attack strategies include enhanced initialization [9], adaptive boundaries [17], multiple target optimization [10], and random restarts with step size scheduling [5]. Further strategies such as utilizing intermediate logits [11], alternating objectives [12], and margin decomposition [7] also contribute to increased robustness. In terms of loss functions, the field has progressed from using cross-entropy to margin loss [3, 7], boundary distance loss [17], DLR [5], and MIFPE loss [13]. Among these, margin loss has shown notable resilience to gradient-related issues, as demonstrated by the Margin Decomposition (MD) attack [7]. The AutoAttack (AA) framework [5] integrates four advanced attack methods—two Auto-PGD (APGD) attacks, a FAB attack [10], and a black-box Square attack [18]—to provide a strong, ensemble-based approach for robust evaluation.
@@ 4.4 Million-Scale Robustness Evaluation
Here, we aim to scale adversarial robustness evaluation up to the million-scale for adversarially trained models on ImageNet. To this end, we first construct a large-scale evaluation dataset named CC1M with 1 million images selected from the CC3M [14] dataset, which is an (image, caption) paired dataset comprising diverse objects, scenes, and visual concepts. We first remove the "unavailable" images, i.e., images showing ‘this image is unavailable’, due to a large number of expired image URLs. Next, we remove the noisy images that do not contain any semantic content, such as random icons. Finally, we apply the Local Intrinsic Dimensionality (LID) [40] to remove anomaly images. The LID is a representation-space metric that has been shown to be able to detect adversarial images [41], backdoor images [42], or low-quality images that are detrimental to self-supervised contrastive learning [43]. Images with deviated LID scores from the average are often viewed as outliers that are geometrically far away from normal data manifold. Therefore, we adopt the Median Absolute Deviation (MAD) with LID as the final score to filter out the outlier (unusual) images. Specifically, we first compute the LID scores for images in CC3M based on their CLIP embeddings, calculate the median of the LID scores, and then select 1 million images with LID scores close to the median.
@@ 7 Introduction
Due to the page limitation of the paper, we further illustrate our method in this supplementary material, which includes the following sections: 1) A detailed analysis of the quantitative results for hyperparameters $K^{\prime}$ and $n$; 2) A comparison of experimental results between the PMA method and the AAA and ACG methods; 3) A comparison of experimental results between the traditional SGD+sign update strategy and optimizer-based strategies; 4) A detailed examination of the ablation results for $P_{max}$ and $P_{y}$ weights; 5) Detailed results of the million-scale adversarial robustness evaluation between the PMA method and other methods.
@@ 8 Detailed quantitative results of hyperparameter $K^{\prime}$ and $n$
For the quantitative investigation of $K^{\prime}$, we compared five different values: 15, 20, 25, 30, and 35. The results, as shown in Table 7, indicate that the optimal performance is achieved when $K^{\prime}$ is set to 25.

### v4

@@ 2 Related Work
The accuracy of robustness evaluations relies heavily on the strength of the attack. White-box attacks used for robustness evaluation must be resilient to issues such as obfuscated gradients [6] and imbalanced gradients [7]. As shown in [6], standard PGD can overestimate robustness in the presence of obfuscated gradients, leading to the development of methods aimed at improving PGD’s effectiveness. These improvements can be broadly classified into two categories: 1) attack strategies and 2) loss functions.
Improved attack strategies include enhanced initialization [9], adaptive boundaries [17], multiple target optimization [10], and random restarts with step size scheduling [5]. Further strategies such as utilizing intermediate logits [11], alternating objectives [12], and margin decomposition [7] also contribute to increased robustness. In terms of loss functions, the field has progressed from using cross-entropy to margin loss [3, 7], boundary distance loss [17], DLR [5], and MIFPE loss [13]. Among these, margin loss has shown notable resilience to gradient-related issues, as demonstrated by the Margin Decomposition (MD) attack [7]. The AutoAttack (AA) framework [5] integrates four advanced attack methods—two Auto-PGD (APGD) attacks, a FAB attack [10], and a black-box Square attack [18]—to provide a strong, ensemble-based approach for robust evaluation.
@@ 4.4 Million-Scale Robustness Evaluation
Here, we aim to scale adversarial robustness evaluation up to the million-scale for adversarially trained models on ImageNet. To this end, we first construct a large-scale evaluation dataset named CC1M with 1 million images selected from the CC3M [14] dataset, which is an (image, caption) paired dataset comprising diverse objects, scenes, and visual concepts. We first remove the "unavailable" images, i.e., images showing ‘this image is unavailable’, due to a large number of expired image URLs. Next, we remove the noisy images that do not contain any semantic content, such as random icons. Finally, we apply the Local Intrinsic Dimensionality (LID) [40] to remove anomaly images. The LID is a representation-space metric that has been shown to be able to detect adversarial images [41], backdoor images [42, 43], or low-quality images that are detrimental to self-supervised contrastive learning [44]. Images with deviated LID scores from the average are often viewed as outliers that are geometrically far away from normal data manifold. Therefore, we adopt the Median Absolute Deviation (MAD) with LID as the final score to filter out the outlier (unusual) images. Specifically, we first compute the LID scores for images in CC3M based on their CLIP embeddings, calculate the median of the LID scores, and then select 1 million images with LID scores close to the median.
@@ 6 Conclusion
Acknowledgements
This work is in part supported by the National Key R&D Program of China (Grant No. 2022ZD0160103) and the National Natural Science Foundation of China (Grant No. 62276067). The computations in this research were performed using the CFFF platform of Fudan University.
@@ 7 Introduction
Due to the page limitation of the paper, we further illustrate our method in this supplementary material, which includes the following sections: 1)Visual illustrations of the attacked images. 2) A detailed analysis of the quantitative results for hyperparameters $K^{\prime}$ and $n$; 3) A comparison of experimental results between the PMA method and the AAA and ACG methods; 4) A comparison of experimental results between the traditional SGD+sign update strategy and optimizer-based strategies; 5) A detailed examination of the ablation results for $P_{max}$ and $P_{y}$ weights; 6) Detailed results of the million-scale adversarial robustness evaluation between the PMA method and other methods. 7) Supplementary experiments on CLIP.
8 isual illustrations of the attacked images.
Figure 2: Visual illustrations of the attacked images.
Fig. 2 visualizes adversarial examples generated by different attacks, along with their predicted labels. We evaluate their quality using PSNR, SSIM, and LPIPS, and report cross-entropy loss between adversarial predictions and ground-truth labels, demonstrating both visual and functional impact.
@@ 9 Detailed quantitative results of hyperparameter $K^{\prime}$ and $n$
For the quantitative investigation of $K^{\prime}$, we compared five different values: 15, 20, 25, 30, and 35. The results, as shown in Table 7, indicate that the optimal performance is achieved when $K$ is set to 25.
@@ 13 Million-Scale adversarial robustness evaluation between the PMA method and other methods
14 Supplementary experiments on CLIP
To address the domain mismatch between Conceptual-Captions and ImageNet, we conducted supplementary experiments using CLIP on CC1M. We tested perturbation ranges of 1, 2, and 3, with a batch size of 32. All samples in a batch—except the text corresponding to the given image—were treated as negative samples. As shown in Table 15, our method consistently achieves the best performance.
$\epsilon$ Clean PGDce PGDdlr PGDmg PGDpm PGDalt PGDmi MD PMA diff 1/255 73.74 26.22 22.32 21.92 21.9 21.4 21.32 21.32 21.16 -0.16 2/255 73.74 13.12 10.45 9.97 9.94 9.48 9.5 9.5 9.4 -0.1 3/255 73.74 7.42 6.2 5.87 5.89 5.5 5.55 5.57 5.45 -0.05
Table 15: Robustness (%) evaluated under varying $\epsilon$.